I need answer for this question
Task 7. Lets assume only books, jeans, and TVs are produced within national economy. 2016 Price, euro Quantity 15 24 425 Year2015 Good Books Jeans TVs Price, euroQuantity 10 27 655 10 If the year 2015 is set as a basic period, calculate for both periods a)Nominal and Real GDP; b) GDP deflator, and explain meaning of the result. c) nflation/Deflation rate; d) Rate of economic growth and explain the result e Calculate tasks a and b, if the year 2016 is set as basic period.
Answer #1

A.

Nominal GDP for 2015 = 10*10+27*6+655*3 = Euro 2227

Since 2015 is the base year, so real GDP for 2015 = Euro 2227

Nominal GDP in 2016 = 15*8 + 24*7 + 425*5 = Euro 2413

Real GDP in 2016 = 10*8 + 27*7 + 655*5 = Euro 3544

B.

GDP deflator = (nominal GDP/real GDP)*100

GDP Deflator in 2015 = (2227/2227)*100 = 100

GDP Deflator in 2016 = (2413/3544)*100 = 68.09

C.

Inflation rate = (68.09-100)/100 = 31.91%

So, there is a deflation rate of 31.91 %.

D.

Real GDP growth rate = (3544-2227)/2227 = 59.14%

So, economic growth rate is positive and it is 59.14%

Pl. repost other unanswered questions for their proper answers!